# Demonstration of logical qubits and repeated error correction with better-than-physical error rates

### v2

@@ Appendix A Statistical analysis and error bars
Appendix B Rejection rates in the Carbon code
@@
Figure 8: Rejection rates during pre-selection (blue circles) and post-selection (orange diamonds) as a function of the number of error correction rounds for the $[[12,2,4]]$ Carbon code. In both cases the expected linear trend (as a function of number of repetitions) is apparent, as is the large separation between pre-rejections and post-rejections.
Appendix C Additional experimental data for the Steane code experiments
The Steane code admits transversal $Y\otimes Y$ measurements, allowing a straight-forward measurement of the process fidelity for the Bell state preparation. For a Bell pair, the average fidelity is given by $F_{a}=(I\otimes I+X\otimes X-Y\otimes Y+Z\otimes Z)/4$, and the process fidelity is $F_{p}=((d+1)F_{a}-1)/d$ [58], where the dimension $d=4$ for the Bell pair. In this section, Table 4, Table 5, and Table 6 provide the process fidelities for the various Steane code experiments and their physical level analogs, and a break-down of the results into the three different Pauli measurements required.
Note that the “gain” values in Table 5 and Table 6 are calculated by dividing the process fidelity (which is calculated using the $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ bases measurements) of the unencoded circuit by the process fidelity of the encoded circuit.
@@ Appendix C Additional experimental data for the Steane code experiments
$E_{p}$ NA $0.69\%_{-0.03\%}^{+0.03\%}$
Table 4: The measured process fidelity $E_{p}$ of the physical-level Bell preparation gadget and the error rates of individual Pauli operators $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$. For comparison, the error rate determined from measuring the probability of getting the wrong parity from $X\otimes X$ and $Z\otimes Z$ is given as $E_{xz}$.
@@
$E_{p}$ NA NA $0.10\%_{-0.04\%}^{+0.06\%}$ $6.8$
Table 5: The measured process fidelity $E_{p}$ of the logical-level Bell preparation gadget utilizing the Steane code where the destructive measurements are analyzed using quantum error correction. The error rates of individual Pauli operators $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ are also reported. For comparison, the error rate determined from measuring the probability of getting the wrong parity from $X\otimes X$ and $Z\otimes Z$ is given as $E_{xz}$.
@@
$E_{p}$ NA NA NA $0.02\%_{-0.01\%}^{+0.03\%}$ $35$
Table 6: The measured process fidelity $E_{p}$ of the logical-level Bell preparation gadget utilizing the Steane code where the destructive measurements are analyzed using QED. The error rates of individual Pauli operators $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ are also reported. For comparison, the error rate determined from measuring the probability of getting the wrong parity from $X\otimes X$ and $Z\otimes Z$ is given as $E_{xz}$.

### v3

@@ Appendix B Statistical analysis and error bars
Appendix C Rejection rates in the Carbon code
@@
Figure 12: Rejection rates during pre-selection (red squares) and post-selection (purple circles) as a function of the number of error correction rounds for the $[[12,2,4]]$ Carbon code. In both cases the expected linear trend (as a function of number of repetitions) is apparent, as is the large separation between pre-rejections and post-rejections.
Appendix D Additional experimental data for the Steane code experiments
The Steane code admits transversal $Y\otimes Y$ measurements, allowing a straight-forward measurement of the state fidelity for the Bell state preparation. For a Bell pair, the state fidelity is given by $F_{s}=(I\otimes I+X\otimes X-Y\otimes Y+Z\otimes Z)/4$. In this section, Table 5, Table 6, and Table 7 provide the state fidelities for the various Steane code experiments and their physical level analogs, and a break-down of the results into the three different Pauli measurements required.
Note that the “gain” values in Table 6 and Table 7 are calculated by dividing the state fidelity (which is calculated using the $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ bases measurements) of the unencoded circuit by the state fidelity of the encoded circuit.
@@ Appendix D Additional experimental data for the Steane code experiments
$E_{s}$ NA $0.70\%_{-0.03\%}^{+0.03\%}$
Table 5: The measured state infidelity $E_{s}=1-F_{s}$ of the physical-level Bell preparation gadget and the error rates of individual Pauli operators $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$. For comparison, the error rate determined from measuring the probability of getting the wrong parity from $X\otimes X$ and $Z\otimes Z$ is given as $E_{xz}$.
@@
$E_{s}$ NA NA $0.10\%_{-0.04\%}^{+0.06\%}$ $6.8$
Table 6: The measured state infidelity $E_{s}=1-F_{s}$ of the logical-level Bell preparation gadget utilizing the Steane code where the destructive measurements are analyzed using quantum error correction. The error rates of individual Pauli operators $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ are also reported. For comparison, the error rate determined from measuring the probability of getting the wrong parity from $X\otimes X$ and $Z\otimes Z$ is given as $E_{xz}$.
@@
$E_{s}$ NA NA NA $0.02\%_{-0.01\%}^{+0.03\%}$ $35$
Table 7: The measured state infidelity $E_{s}=1-F_{s}$ of the logical-level Bell preparation gadget utilizing the Steane code where the destructive measurements are analyzed using QED. The error rates of individual Pauli operators $X\otimes X$, $Y\otimes Y$, and $Z\otimes Z$ are also reported. For comparison, the error rate determined from measuring the probability of getting the wrong parity from $X\otimes X$ and $Z\otimes Z$ is given as $E_{xz}$.